If a plane is flying at an altitude of 500m and the bottle lands 400m horizontally from the initial dropping point, how fast was the plane flying when the bottle was released.
-Physics-

Answer by RAY100(1637)

The vertical drop is only related to the height and gravity by,
.
d= 1/2 gt^2,,,,where d=distance (m),,,g=9.81 (m / sec^2),,, and t=time (sec)
.
500 m = 1/2 *9.81 *t^2
.
500*2/9.81 = 101.936 sec^2 = t^2
.
10.0964 sec = t,,,,
.
.
the velocity of the plane in that time is distance / time
.
v=d/t = 400/10.0964 = 39.618 m/sec
.